OpenStudy (anonymous):

PLEASE HELP! Which equation can be used to solve this problem? Juanita hiked 5 1/4 miles on Saturday. That was 3/4 as far as Laura hiked. How far did Laura hike? A. m= 3/4 times 5 1/4 B. 3/4m=5 1/4 C. 3/4 divided by m=5 1/4 D. m minus 3/4=5 1/4

OpenStudy (anonymous):

help? @Crissy15

OpenStudy (texaschic101):

miles Laura hiked is represented by " m " 3/4m = 5 1/4 3/4m = 21/4 m = (21/4) / (3/4) m = 21/4 * 4/3 m = 7 check.. 3/4(7) = 21/4 21/4 = 21/4 (correct) So Laura hiked 7 miles and Juanita hiked 3/4 of 7 miles which equals 5 1/4 miles

OpenStudy (anonymous):

so the answer is B? @texaschic101

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

yes...answer is B